1. Richard has available 400 yards of fencing and wishes to enclose a rectangular area.
a.) Express the area A of the rectangle as a function of the width x of the rectangle?
b.) What is the domain of A?

Found 3 solutions by ikleyn, Theo, Solver92311:
Answer by ikleyn(52795)

.
Richard has available 400 yards of fencing and wishes to enclose a rectangular area.
a.) Express the area A of the rectangle as a function of the width x of the rectangle?
b.) What is the domain of A?
~~~~~~~~~~~~~~~~~~

(a)  the Length + the Width = 400/2 = 200,

     so if the length is x yards, then the width is (200-x) yards.


     Then the area is A = xy square yards, which is A(x) = x*(200-x) = -x^2 + 200x.



(b)  The function is A(x) = x*(200-x) = -x^2 + 200x,

     and it makes sense where A(x) is non-negative and x is non-negative, 

     so the domain of the function A(x), defined in this way for the area, is  0 <= x <= 200.

Solved.



Answer by Theo(13342)

a = l * w
p = 2 * (l + w)
a is the area
l is the length
w is the width
p is the perimeter

when p = 400, p = 2 * (l + w) becomes 400 = 2 * (l + w)
divide both sides of this equation by 2 to get:
200 = l + w
solve for l to get:
l = 200 - w

in the formula for area:
when l = 200 - 2, a = l * w becomes:
a = (200 - w) * w.
simplify to get:
a = 200 * w - w^2.

a must be greater than or equal to 0 (can't be negative).
if a >= 0, then:
200 * w - w^2 >= 0
add w^2 to both sides of this equation to get:
200 * w >= w^2
divide both sides of this equation by w to get:
200 >= w

if 200 >= w, then:
w <= 200
w must also be >= 0, therefore:
0 <= w <= 200

that's your domain.
the area is your range.

the domain is the width which is greater than or equal to 0 and less than or equal to 200.
that says the domain of the area is 0 <= width <= 200.
that should be your solution.

the equation of area = 200 * width minus width squared can be graph by letting y = the area and x = the width.
the formula for graphing becomes y = 200x - x^2.

the graph looks like this:



you can see from the grpah that the area, represented by y, is not negative when 0 <= x <= 200, with x representing the width.

the coordinate points on the graph are in (x,y) format.
x represents the width of the rectangle.
y represents the area of the rectangle.
at (100,10,000) the width is 100 and the area is 10,000.
you don't see the length on the graph, but you can always find the length, because length + width will always be 200.
therefore, when the width is 100, the length has to be 100.
the area is equal to length * width = 100 * 100 = 10,000, as shown on the graph.

Answer by Solver92311(821)



The perimeter of a rectangle is given by

So the length for a fixed perimeter as a function of the width is

Then since the area is length times width, the area as a function of the width with a fixed perimeter is .

Given a perimeter of 400 yards and a width of , the function you seek is:



Given that an area less than zero is absurd, the domain of is the set of all such that is non-negative, namely the closed interval between the two zeros of the function. I leave it as an exercise for the student to find the endpoints of the domain interval.

John

My calculator said it, I believe it, that settles it

From
I > Ø